integrate 1/1+x^4

Grade:11

1 Answers

Shyama P
12 years ago

1/1+x4 =1/[x2(x2+1/x2)]

Split as :

1+1/x2 - (1-1/x2)

    2(x2+1/x2)

Evaluate as 2 integrals :

Put t=x-1/x in 1st & t=x+1/x in 2nd